OpenStudy (anonymous):

. In a test run, a certain car accelerates uniformly from zero to 24.0 m/s in 2.95 s. (a) What is the magnitude of the car’s acceleration? (b) How long does it take the car to change its speed from 10.0 m/s to 20.0 m/s? (c) Will doubling the time always double the change in speed? Why?

OpenStudy (cwrw238):

(a) acceleration = velocity/ time = 24/2.95 m/s/s

OpenStudy (anonymous):

8.136 m/s^2

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

b) 20 - 10 ------ = 8.136 t solve for t

OpenStudy (anonymous):

t = 1.229

OpenStudy (cwrw238):

yea

OpenStudy (cwrw238):

(c) acceleration = ( v 2 - v1)/ t = 2(v2-v1) / 2t yes - because acceleration is constant